Marginal cost is defined as: the change in total
Answer #1

Answer A is correct.

Definition of Marginal cost is -: Marginal cost is the change in total cost from producing one more unit output.

Answer A is correct.

The marginal cost curve often decrease at first and then statrts to increase. this is explained by the law of diminishing returns.

Marginal cost for 3rd unit is $20. where n=3 TCn = 65 , TC(n-1) = 45  

MCn = TCn - TC(n-1) , using this formula

= 65 - 45

MCn = $20